• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Lindgren

    Log entry time 07:13:19 on October 7,2002

    Entry number 87551

    keyword=shift summary

    00:05 Am Continuing run 1191
    00:06 End run 1191 1M events, Start run 1192
    00:38 End Run 1192, 1M events, Start run 1193
    01:11 End run 1193 1M events, start run 1194
    01:15 Helium level in Q1 continues to vary between 35 and 60 over a 5 minute interval
    01:45 End run 1194 1M events, start run 1195
    01:50 Helium level varies between 60 and 70 over a 5 minute interval
    01:58 TS1 crashed
    02:00 Ended run 1195 337K events, start run 1196
    02:35 Ended run 1196 1M events, start run 1197
    03:08 Ended run 1197 1M events, start run 1198
    03:40 Ended run 1198 1m events, start run 1199
    04:12 Ended run 1199 1M events, start run 1200
    04:30 MCC called to do a QE run. Will take 10 minutes, Ending run 1200.
    04:35 MCC called to cancel their measurement. Start run 1201
    05:08 Ended run 1201 1M events, start run 1202
    05:42 Ended run 1202 1M events, start run 1203
    05:54 Ended run 1203 332K events stopped run to warm up target
    06:05 Both left and right Q1 crashed
    JP Chen came in at 6:00 am to warm up the target.
    Stopped beam at 6:00 am.
    Took elastic data from midnight to 6:00 am. Q1 and Q2 and the right dipole crashed at 6:00 am. The dipole came back after resetting it.
    We did not reset Q1 and Q2.